Write the equation of a line with a slope of 2 that passes through the point (3, 7).
Answer by mananth(16949)

(3,7) slope 2
equation of line =
y-y1=m(x-x1) where (x1,y1) is the point on the line
y-7=2(x-3)
y-7=2x-6
y-2x=1
y-2x=1 is the required equation